# PyMieDiff: A differentiable Mie scattering library

### v2

@@ Acknowledgements.
The near-fields can be found by matching the continuity conditions at each spherical interface, which yields following formulae:Bohren and Huffman (1998)
$\mathbf{E}_{i}=\sum_{n=1}^{\infty}E_{n}\left(\mathbf{M}_{o1n}^{(1)}-i\mathbf{N}_{e1n}^{(1)}\right),$
$\mathbf{H}_{i}=\frac{-k}{\omega\mu}\sum_{n=1}^{\infty}E_{n}\left(\mathbf{M}_{e1n}^{(1)}+i\mathbf{N}_{o1n}^{(1)}\right),$
$\mathbf{E}_{1}=\sum_{n=1}^{\infty}E_{n}\left(c_{n}\mathbf{M}_{o1n}^{(1)}-id_{n}\mathbf{N}_{e1n}^{(1)}\right),$
$\mathbf{H}_{1}=\frac{-k_{1}}{\omega\mu_{1}}\sum_{n=1}^{\infty}E_{n}\left(d_{n}\mathbf{M}_{e1n}^{(1)}+ic_{n}\mathbf{N}_{o1n}^{(1)}\right),$
$\mathbf{E}_{2}=\sum_{n=1}^{\infty}E_{n}\left(f_{n}\mathbf{M}_{o1n}^{(1)}-ig_{n}\mathbf{N}_{e1n}^{(1)}+v_{n}\mathbf{M}_{o1n}^{(2)}-iw_{n}\mathbf{N}_{e1n}^{(2)}\right),$
$\mathbf{H}_{2}=\frac{-k_{2}}{\omega\mu_{2}}\sum_{n=1}^{\infty}E_{n}\left(g_{n}\mathbf{M}_{e1n}^{(1)}+if_{n}\mathbf{N}_{o1n}^{(1)}+w_{n}\mathbf{M}_{e1n}^{(2)}+iv_{n}\mathbf{N}_{o1n}^{(2)}\right),$
$\mathbf{E}_{3}=\sum_{n=1}^{\infty}E_{n}\left(ia_{n}\mathbf{N}_{e1n}^{(3)}-b_{n}\mathbf{M}_{o1n}^{(3)}\right),$
$\mathbf{H}_{3}=\frac{k}{\omega\mu}\sum_{n=1}^{\infty}E_{n}\left(ib_{n}\mathbf{N}_{o1n}^{(3)}+a_{n}\mathbf{M}_{e1n}^{(3)}\right).$
The subscripts indicate the region and type of field. “i”: incident field in the background medium, “1”: total field in the core, “2”: total field in the shell medium, and “3”: scattered field in the background medium. The total fields in the background medium are given by $\mathbf{E}_{\text{tot}}=\mathbf{E}_{i}+\mathbf{E}_{3}$ and $\mathbf{H}_{\text{tot}}=\mathbf{H}_{i}+\mathbf{H}_{3}$.
.9 Second derivatives of the Spherical Bessel functions

### v3

@@ Acknowledgements.
The near-fields in the $l$th region of the sphere can be found by matching the continuity conditions at each spherical interface, which yields following formulae:Ladutenko et al. (2017)
$\mathbf{E}_{l}=\sum_{n=1}^{\infty}E_{n}\left[c_{n}^{(l)}\mathbf{M}_{o1n}^{(1)}-id_{n}^{(l)}\mathbf{N}_{e1n}^{(1)}+ia_{n}^{(l)}\mathbf{N}_{e1n}^{(3)}-b_{n}^{(l)}\mathbf{M}_{o1n}^{(3)}\right],$ (39)
$\mathbf{H}_{l}=\frac{k_{l}}{\omega\mu}\sum_{n=1}^{\infty}E_{n}\left[d_{n}^{(l)}\mathbf{M}_{e1n}^{(1)}+ic_{n}^{(l)}\mathbf{N}_{o1n}^{(1)}-ib_{n}^{(l)}\mathbf{N}_{o1n}^{(3)}-a_{n}^{(l)}\mathbf{M}_{e1n}^{(3)}\right],$ (40)
where $a_{n}^{(l)}$, $b_{n}^{(l)}$, $c_{n}^{(l)}$ and $d_{n}^{(l)}$ are the expansion coefficients. These have been found by Ladutenko, Konstantin, et al. Ladutenko et al. (2017) to be,
$a_{n}^{(l)}=\frac{D_{n}^{(1)}\left(m_{l}x_{l}\right)T_{1}\left(m_{l+1}x_{l}\right)+T_{3}\left(m_{l+1}x_{l}\right)m_{l}/m_{l+1}}{\xi_{n}\left(m_{l}x_{l}\right)U\left(m_{l}x_{l}\right)},$ (41)
$b_{n}^{(l)}=\frac{D_{n}^{(1)}\left(m_{l}x_{l}\right)T_{2}\left(m_{l+1}x_{l}\right)m_{l}/m_{l+1}+T_{4}\left(m_{l+1}x_{l}\right)}{\xi_{n}\left(m_{l}x_{l}\right)U\left(m_{l}x_{l}\right)},$ (42)
$c_{n}^{(l)}=\frac{D_{n}^{(3)}\left(m_{l}x_{l}\right)T_{2}\left(m_{l+1}x_{l}\right)m_{l}/m_{l+1}+T_{4}\left(m_{l+1}x_{l}\right)}{\psi_{n}\left(m_{l}x_{l}\right)U\left(m_{l}x_{l}\right)},$ (43)
$d_{n}^{(l)}=\frac{D_{n}^{(3)}\left(m_{l}x_{l}\right)T_{1}\left(m_{l+1}x_{l}\right)+T_{3}\left(m_{l+1}x_{l}\right)m_{l}/m_{l+1}}{\psi_{n}\left(m_{l}x_{l}\right)U\left(m_{l}x_{l}\right)},$ (44)
where,
$U(z)=D_{n}^{(1)}(z)-D_{n}^{(3)}(z),$ (45)
$T_{1}(z)=a_{n}^{(l+1)}\xi_{n}(z)-d_{n}^{(l+1)}\psi_{n}(z),$ (46)
$T_{2}(z)=b_{n}^{(l+1)}\xi_{n}(z)-c_{n}^{(l+1)}\psi_{n}(z),$ (47)
$T_{3}(z)=d_{n}^{(l+1)}D_{n}^{(1)}(z)\psi_{n}(z)-a_{n}^{(l+1)}D_{n}^{(3)}(z)\xi_{n}(z),$ (48)
$T_{4}(z)=c_{n}^{(l+1)}D_{n}^{(1)}(z)\psi_{n}(z)-b_{n}^{(l+1)}D_{n}^{(3)}(z)\xi_{n}(z).$ (49)
The total fields in the background medium are given by $\mathbf{E}_{\text{tot}}=\mathbf{E}_{i}+\mathbf{E}_{s}$ and $\mathbf{H}_{\text{tot}}=\mathbf{H}_{i}+\mathbf{H}_{s}$, where,
$\mathbf{E}_{i}=\sum_{n=1}^{\infty}E_{n}\left(\mathbf{M}_{o1n}^{(1)}-i\mathbf{N}_{e1n}^{(1)}\right),$ (50)
$\mathbf{H}_{i}=\frac{k_{l}}{\omega\mu}\sum_{n=1}^{\infty}E_{n}\left(\mathbf{M}_{e1n}^{(1)}+i\mathbf{N}_{o1n}^{(1)}\right),$ (51)
$\mathbf{E}_{s}=\sum_{n=1}^{\infty}E_{n}\left(ia_{n}\mathbf{N}_{e1n}^{(3)}-b_{n}\mathbf{M}_{o1n}^{(3)}\right),$ (52)
$\mathbf{H}_{s}=\frac{k_{l}}{\omega\mu}\sum_{n=1}^{\infty}E_{n}\left(-ib_{n}\mathbf{N}_{o1n}^{(3)}-a_{n}\mathbf{M}_{e1n}^{(3)}\right).$ (53)
Physically, to ensure that the electromagnetic fields remain finite at the origin of the sphere, the corresponding expansion coefficients must vanishBohren and Huffman (1998), requiring $a_{n}^{(1)}=b_{n}^{(1)}=0$. Furthermore, it can be shown Ladutenko et al. (2017) that, an inspection of Eqs. (39), (40) and Eqs. (50)-(53), reveals that $c_{n}^{L+1}=d_{n}^{L+1}=1$.
.9 Spherical bessel based implementation for testing
During development, we also implemented SciPy wrappers and PyTorch-based recurrences for standard spherical Bessel functions ($h^{1}_{n}(z)$, $j_{n}(z)$ and $y_{n}(z)$) to test a core-shell Mie formulation. Although computing Mie scattering directly with these functions leads to numerical instability for large or absorbing particles, providing these Bessel functions in PyTorch may be valuable for researchers tackling other physical systems.
$h^{1}_{n}(z)=j_{n}(z)+\text{i}\,y_{n}(z)\,.$ (54)
$f_{n}^{\prime}(z)$ $=f_{n-1}(z)-\frac{n+1}{z}f_{n}(z)\,$, (55)
$f_{n}^{\prime}(z)$ $=-f_{n+1}(z)+\frac{n}{z}f_{n}(z)\,$,
where $f_{n}(z)$ is any spherical bessel function. Note: In practice, to avoid negative orders, derivatives for $n>0$ are calculated using the first relation in Eqs. (55), while the derivative for $n=0$ is obtained from the second relation:
$f_{0}^{\prime}(z)\;=\;-f_{1}(z)\,.$ (56)